# Evil twins are not that evil: Qualitative insights into machine-generated prompts

### v3

@@ Appendix B Token replacement examples
We show randomly picked examples of single-token autoprompt replacements that do not affect the continuation, have a moderate effect on it or have a strong effect on it in tables 10, 11 and 12, respectively. In addition, Table B provides additional examples of replacement with null-effect, where tokens are colored based on the number of equivalent substitutes.
@@
[…]
@@
Table 13: Additional Examples of replacement for autoprompts and original prompts (10 last tokens). Color represents the number of substitutes: $>50$ substitutes and $<50$ substitutes. When there are less than 50 substitutes, a random subset is displayed. Difficult-to-render characters are replaced by ’?’.
@@ Appendix E Assets
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
@@
Table 17: Pythia-1.3B: reproducing pruning and replacement experiments.
@@
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
@@
Table 18: OLMo-1B: reproducing pruning and replacement experiments.
@@
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
@@
Table 19: OLMo-7B: reproducing pruning and replacement experiments.
@@
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
@@
Table 20: Pythia-12B: reproducing pruning and replacement experiments.
@@
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
@@
Table 21: OLMo-7B-Instruct: reproducing pruning and replacement experiments.

### v4

@@ Appendix B Token replacement examples
We show randomly picked examples of single-token autoprompt replacements that do not affect the continuation, have a moderate effect on it or have a strong effect on it in tables 9, 10 and 11, respectively. In addition, Table 12 provides additional examples of replacement with null-effect, where tokens are colored based on the number of equivalent substitutes.
gray!50
@@
ipzig: ’ century’, ’ society’, ’ secret’, ’ glad’, ’ revolution’, ’ intellectual’, ’ consciousness’, …
@@
Table 12: Additional Examples of replacement for autoprompts and original prompts (10 last tokens). Color represents the number of substitutes: >50>50 substitutes and <50<50 substitutes. When there are less than 50 substitutes, a random subset is displayed. Difficult-to-render characters are replaced by ’?’.
black
@@
Semantic consistency of the equivalent sets
Table 5 provides the semantic similarity between substitutes.
(a) Autoprompts
(b) Original Prompts
@@ Appendix E Assets
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original prompt).
@@
Figure 6: Pythia-1.3B: Reproducing pruning and replacement experiments.
@@
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original prompt).
@@
Figure 7: OLMo-1B: Reproducing pruning and replacement experiments.
@@
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original prompt).
@@
Figure 8: OLMo-7B: Reproducing pruning and replacement experiments.
@@
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original prompt).
@@
Figure 9: Pythia-12B: Reproducing pruning and replacement experiments.
@@
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original prompt).
@@
Figure 10: OLMo-7B-Instruct: Reproducing pruning and replacement experiments.